OpenStudy (anonymous):

The length of a rectangle is 7 units more than its width. If the width is doubled and the length is increased by 2, the area is increased by 42 square units. The width is how many units and the length is how many units?

OpenStudy (anonymous):

l = 7+w lw+42 = 2w*(l+2) lw + 42 = 2wl + 2w wl +2w = 42 w(7+w) + 2w - 42 = 0 w^2 + 9w -42 = 0 Then solve for w & l :D

OpenStudy (anonymous):

ehhhh..... just tell me the answer please im odee tired :)